# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

### v1

@@ 4.4 Pendulum under CBRL Piecewise-LQR
Recalling that the LQR controller is not sufficient to solve the Pendulum problem, we consider a piecewise-LQR controller that partitions the state space into four regions (see Fig. 11 in Appendix C.4). The target state $x^{*}=[\theta^{*},\dot{\theta}^{*}]$ is selected based on the boundary angle in each partition (counter-clockwise boundary angle if $\dot{\theta}>0$, and clockwise boundary angle otherwise), and the energy conservation law for target angular velocity (see Appendix C.4 for more details). We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics (13) (Appendix C.4), solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned (Fig. 4(c)).
Fig. 4 presents numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds. All parameters in (13) are initialized uniformly within $(0,1)$ and then learned using our control-policy-parameter gradient ascent iteration (5). Fig. 4(a) illustrates a comparable performance between our CBRL approach and the best baseline DDPG in terms of the convergence w.r.t. the number of episodes, while Fig. 4(b) clearly demonstrates that CBRL outperforms all other baselines w.r.t. running time, in addition to demonstrating a more stable training process.
5 Discussion and Limitations
In this paper we devise a CBRL approach to directly learn the optimal policy, establishing its theoretical properties, adapting it to the LQR control-theoretic framework, and empirically demonstrating its performance benefits. We now turn to discuss limitations of our CBRL approach. Although primarily used for theoretical purposes, the supremum of our CBRL operator (2) and its fixed point (3) may be challenging if computed directly when the parameter space is continuous. This issue, however, is no different than corresponding challenges in standard RL when the action space is continuous, and thus similar steps can be taken to address the issue which does not arise in our Algorithm 1. The stepsize selection and parameter-vector initialization are also important implementation issues, just like in nonconvex optimization and RL. As described in Appendix C, the former issue can be addressed with similar steps taken in standard RL, including the adaptive stepsize selection in our implementation of Algorithm 1, and experimental results demonstrate the robustness of our CBRL approach w.r.t. the latter issue. Given the efficiency of our CBRL approach relative to state-of-the-art methods, additional computations for stepsize selection and parameter-vector initialization can be employed when issues are encountered while retaining overall computational efficiencies. Future work includes exploring other control-theoretic frameworks that can solve problems with different characteristics. For instance, consideration of Model Predictive Control as the control-theoretic framework instead of LQR would allow us to impose state constraints related to, e.g., safety.
Acknowledgments
This work was supported by IBM through the IBM-Rensselaer Future of Computing Research Collaboration.
@@ Appendix A Control-Based Reinforcement Learning Approach
In this appendix, we present the proofs of our main theoretical results on control-based RL together with additional results and technical details related to Section 2. We also present the algorithmic details of our control-policy-parameter gradient ascent method used for the results in Section 4.
A.1 Proof of Lemma 2.1.
Lemma 2.1. For any $\gamma\in(0,1)$, operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm. Suppose Assumption 2.1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same optimal outcome as the Bellman operator.
@@
For the operator ${{\mathbf{T}}}$ defined on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ in (2) and for any two functions $\tilde{q}_{1}(x,{\mathcal{F}}_{p_{1}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ and $\tilde{q}_{2}(x,{\mathcal{F}}_{p_{2}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with $p_{1},p_{2}\in{{\mathbb{S}}}$, we obtain
$\|{{\mathbf{T}}}\tilde{q}_{1}-{{\mathbf{T}}}\tilde{q}_{2}\|_{\infty}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-r(x,{\mathcal{F}}_{p}(x))-\gamma\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{|}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{|}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\sup_{z\in{{\mathbb{X}}},p^{\prime}\in{{\mathbb{S}}}}\left|\tilde{q}_{1}(z,{\mathcal{F}}_{p^{\prime}})-\tilde{q}_{2}(z,{\mathcal{F}}_{p^{\prime}})\right|$
$\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\|\tilde{q}_{1}-\tilde{q}_{2}\|_{\infty}$
@@
where (a) is by definition in (2), (b) follows from straightforward algebra, (c) and (d) are due to the triangle inequality, and (e) and (f) directly follow by definition. For any $\gamma\in(0,1)$, this establishes that the operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm, thus rendering the desired result for the first part.
@@
For the second part of the lemma, under the stated supposition, we know that the optimal policy $\pi^{\ast}(x)$ realized by the Bellman operator holds for a unique vector $p^{\prime}$ in the parameter set ${{\mathbb{S}}}$ and any $x\in{{\mathbb{X}}}$. We also know that the Bellman operator and the operator ${{\mathbf{T}}}$ in (2) are both contractions in supremum norm with unique fixed points, where the fixed point of the Bellman operator is the optimal solution of the Bellman equation. By showing that these two fixed points coincide, the desired result follows.
To this end, from the definition of $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})$ and under the stated supposition, we have $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})=q(x,{\mathcal{F}}_{p^{\prime}}(x))=q(x,\pi^{\ast}(x))$ for any $x\in{{\mathbb{X}}}$, where $p^{\prime}$ is the unique vector in the parameter set ${{\mathbb{S}}}$ associated with $\pi^{\ast}(x)$. Then, from the right hand side of (2), we conclude $\tilde{q}(x,{\mathcal{F}}_{p^{\prime}})=q(x,{\mathcal{F}}_{p^{\prime}}(x))=q(x,\pi^{\ast}(x))\leq\sup_{p^{{}^{\prime\prime}}\in{{\mathbb{S}}}}\tilde{q}(x,{\mathcal{F}}_{p^{\prime\prime}})$. However, the last inequality must hold with equality due to the optimality of the policy $\pi^{\ast}(x)$, and therefore the fixed point of the operator ${{\mathbf{T}}}$ coincides with the fixed point of the Bellman operator, thus rendering the desired result. ∎
A.2 Proof of Theorem 2.1.
@@
From Lemma 2.1, for $i\in[2]$, we have that the contraction operators ${{\mathbf{T}}}_{i}:{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})$ under the parameter sets ${{\mathbb{S}}}_{i}$ converge to the unique fixed points $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ which satisfy, for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$,
$\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}^{(i)}_{p})=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}^{(i)}_{p}(x)}(x,y)\bigg{[}r(x,{\mathcal{F}}^{(i)}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}_{i}}\tilde{q}_{i}^{\prime}(y,{\mathcal{F}}^{(i)}_{p^{\prime}})\bigg{]}.$ (8)
@@
We first introduce the following lemma used in the proof of Theorem 2.1.
Lemma A.1.

### v2

@@ Introduction
Recall that the LQR controller is not sufficient to solve the Pendulum problem, even if all the parameters of the system are known (e.g., mass of the link $m$, length of the link $l$, moment of inertia of the link $J$, and gravity $g$), because a nonlinear controller is required. Consequently, we consider a piecewise-LQR controller that partitions the state space into four regions (LQR 1 – 4 in Fig. 11 of the appendix). In terms of the target state $x^{*}=[\theta^{*},\dot{\theta}^{*}]$, the angle $\theta^{*}$ is selected based on the boundary angle in each partition (counter-clockwise boundary angle if $\dot{\theta}>0$, and clockwise boundary angle otherwise), while the angular velocity $\dot{\theta}^{*}$ is selected based on the energy conservation law; see the appendix for more details. We address the problem within the context of our CBRL approach by exploiting the general matrix form for the piecewise-LQR dynamics given by (18) in the appendix, solely taking into account general physical relationships and laws, with the unknown parameters $a_{0},a_{1},b_{0},c_{0},d_{0}$ to be learned; refer to Fig. 4(c).
Fig. 4 and Table 7 (see appendix) present numerical results for the three state-of-the-art baselines (continuous actions) and our CBRL approach, with each run over five independent random seeds; see Table 8 and Fig. 12 in the appendix for the four sets of initial parameters in (18). Fig. 4(a) – (b), Table 7 and Fig. 12(a) – (b) clearly demonstrate that our CBRL approach provides far superior performance (both mean and std. dev.) over all baselines w.r.t. both the number of episodes and running time upon convergence of our CBRL algorithm after a relatively small number of episodes. In particular, even with only four partitions for the difficult nonlinear Pendulum RL task, our CBRL approach provides significantly better mean performance, together with a much lower std. dev., than the closest competitor DDPG after $150$ episodes. These numerical results also demonstrate a more stable training process for our CBRL approach. Finally, Fig. 12(c) – 12(f) illustrates the learning behavior of CBRL parameters given different initialization.
Discussion
Our results demonstrate an important form of robustness exhibited by the optimal control policy of our CBRL approach w.r.t. the learned parameter values. In particular, we performed additional comparative numerical experiments to evaluate the performance of LQR with the known true parameters for Cart Pole and the performance of piecewise-LQR with the known true parameters for Mountain Car and Pendulum (Lunar Lander is omitted because its underlying true parameters are not known to us). We then compare the relative difference between these results for LQR with the true parameters and the corresponding results of our CBRL approach in Figure 1(a),(b), Figure 3(a),(b), and Figure 4(a),(b). It is important to note that the parameter values learned by our CBRL approach and presented in Figure 1(c), Figure 3(c), and Figure 4(c) differ considerably from the corresponding true parameters. Despite these non-negligible parameter value differences, the comparative relative performance differences for Cart Pole and Pendulum respectively show that LQR and piecewise-LQR with the true parameters provide no improvement in return over the corresponding return of our CBRL approach in Figure 1(a),(b) and Figure 4 (a),(b). In addition, the return of our CBRL approach for Mountain Car in Figure 3(a),(b) is within $2\%$ of the corresponding return of piecewise-LQR with the true parameters.
The supremum over the parameter space in the definition of our CBRL operator (CBRL Approach) and its unique fixed point (4) are primarily used for theoretical purposes in this paper. However, one might be concerned about potential computational challenges associated with the supremum when the parameter space is continuous. We first note that this issue is not fundamentally different from the corresponding challenges in standard RL when the action space is continuous, and thus similar steps can be taken to address the issue. Another important factor that mitigates such concerns is the aforementioned robustness of the optimal control policy w.r.t. the learned parameter values. Lastly, we note that this issue did not arise in our numerical experiments using Algorithm 1.
One important implementation issue is stepsize selection, just like in nonconvex optimization and RL. We address this issue by using similar methods employed in standard RL, i.e., the adaptive stepsize selection in our implementation of Algorithm 1. Specifically, we reduce the stepsize $\eta$ in (6) by a factor of $0.99$ each time achieving the “solved” return. Another important implementation issue concerns parameter-vector initialization, again like in nonconvex optimization and RL. As discussed for each of the above RL tasks, the experimental results in Figs. 681012 demonstrate the robustness of our CBRL approach w.r.t. parameter-vector initialization. An additional important factor here is the aforementioned robustness of the optimal control policy w.r.t. the learned parameter values. Moreover, given the efficiency of our CBRL approach relative to state-of-the-art methods, additional computations for stepsize selection and parameter-vector initialization can be employed when issues are encountered while retaining overall computational efficiencies.
Conclusion
In this paper we devise a CBRL approach to support direct learning of the optimal policy. We establish various theoretical properties of our approach, including convergence and optimality of our CBRL operator, a new control-policy-parameter gradient ascent theorem, and a gradient ascent algorithm based on this theorem. We then adapt our CBRL approach to the LQR control-theoretic framework, as a representative example, and evaluate its performance on several classical RL tasks. These empirical results demonstrate the significant benefits of our CBRL approach over state-of-the-art baseline methods in terms of improved quality and robustness of the solution and reduced sample complexity and running time.
@@ Appendix A Control-Based Reinforcement Learning Approach
In this appendix, we present the proofs of our main theoretical results on control-based RL together with additional results and technical details related to our CBRL approach of the main body of the paper. We also present the algorithmic details of our control-policy-parameter gradient ascent method used for our experimental results of the main body of the paper and in the appendix below.
Proof of Theorem 1.
Assumption 1. There exist a policy function ${\mathcal{G}}$ in the family ${\mathbb{F}}$ and a unique vector $p^{\ast}$ in the parameter set ${{\mathbb{S}}}$ such that, for any state $x\in{{\mathbb{X}}}$, $\pi^{\ast}(x)={\mathcal{F}}_{p^{\ast}}(x)={\mathcal{G}}(p^{\ast})(x)$.
Theorem 1. (a) For any $\gamma\in(0,1)$, the operator ${{\mathbf{T}}}$ in (CBRL Approach) is a contraction in the supremum norm. (b) Suppose Assumption 1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same asymptotically optimal outcome as that of the Bellman operator ${{\mathcal{T}}_{B}}$.
@@
For the operator ${{\mathbf{T}}}$ defined on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ in (CBRL Approach) and for any two functions $\tilde{q}_{1}(x,{\mathcal{F}}_{p_{1}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ and $\tilde{q}_{2}(x,{\mathcal{F}}_{p_{2}})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with $p_{1},p_{2}\in{{\mathbb{S}}}$, we obtain
$\|{{\mathbf{T}}}\tilde{q}_{1}-{{\mathbf{T}}}\tilde{q}_{2}\|_{\infty}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-r(x,{\mathcal{F}}_{p}(x))-\gamma\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\Big{|}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\Big{[}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{]}\Big{|}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\Big{|}\sup_{p_{1}\in{{\mathbb{S}}}}\tilde{q}_{1}(y,{\mathcal{F}}_{p_{1}})-\sup_{p_{2}\in{{\mathbb{S}}}}\tilde{q}_{2}(y,{\mathcal{F}}_{p_{2}})\Big{|}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\sup_{z\in{{\mathbb{X}}},p^{\prime}\in{{\mathbb{S}}}}\left|\tilde{q}_{1}(z,{\mathcal{F}}_{p^{\prime}})-\tilde{q}_{2}(z,{\mathcal{F}}_{p^{\prime}})\right|$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\sup_{x\in{{\mathbb{X}}},p\in{{\mathbb{S}}}}\gamma\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\|\tilde{q}_{1}-\tilde{q}_{2}\|_{\infty}$
@@
where (a) is by definition in (CBRL Approach), (b) follows from straightforward algebra, (c) and (d) are due to the triangle inequality, and (e) and (f) directly follow by definition. For any $\gamma\in(0,1)$, this establishes that the operator ${{\mathbf{T}}}$ in (CBRL Approach) is a contraction in the supremum norm, thus rendering the desired result for the first part of the theorem.
@@
For the second part of the theorem, under the stated supposition, we know that the optimal policy $\pi^{\ast}(x)$ realized by the Bellman operator holds for a unique vector $p^{\ast}$ in the parameter set ${{\mathbb{S}}}$ and any $x\in{{\mathbb{X}}}$. We also know that the Bellman operator ${{\mathcal{T}}_{B}}$ in (2) and our CBRL operator ${{\mathbf{T}}}$ in (CBRL Approach) are both contractions in supremum norm with unique fixed points, where the fixed point of the Bellman operator is the optimal solution of the Bellman equation. In particular, repeatedly applying the Bellman operator ${{\mathcal{T}}_{B}}$ in (2) to an action-value function $q(x,u)\in{\mathbb{Q}}({{\mathbb{X}}}\times{\mathbb{U}})$ is well-known to asymptotically yield the unique fixed point equation ${{\mathcal{T}}_{B}}q_{B}^{\ast}=q_{B}^{\ast}$ where
$q_{B}^{\ast}(x,\pi^{\ast}(x))$ $=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,\pi^{\ast})\Big{[}r(x,\pi^{\ast}(x))+\gamma\sup_{u^{\prime}\in{\mathbb{U}}}q_{B}^{\ast}(y,u^{\prime})\Big{]}=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,\pi^{\ast})\Big{[}r(x,\pi^{\ast}(x))+\gamma q_{B}^{\ast}(y,\pi^{\ast}(y))\Big{]},$ (9)
the second equality follows from the definition of $\pi^{\ast}(x)$, and $q_{B}^{\ast}$ is the optimal action-value function (Szepesvari 2010). Likewise, repeatedly applying our CBRL operator ${{\mathbf{T}}}$ in (CBRL Approach) to an action-value function $\tilde{q}(x,{\mathcal{F}}_{p})\in\tilde{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi)$ asymptotically renders the unique fixed point equation ${{\mathbf{T}}}\tilde{q}_{C}^{\ast}=\tilde{q}_{C}{{}^{\ast}}$ where
$\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ $=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma\tilde{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\ast}})\Big{]},$ (10)
and the second equality follows from the relationship between $\tilde{q}(x,{\mathcal{F}}_{p^{\ast}})$ and $\tilde{q}(x,\pi^{\ast})$, given below in (11), under the supposition of the theorem together with the optimality of $\pi^{\ast}$. By showing that these two fixed points $q_{B}^{\ast}(x,\pi^{\ast}(x))$ and $\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ in (9) and (10) coincide, the desired result follows.
To this end, from the definition of $\tilde{q}(x,{\mathcal{F}}_{p^{\ast}})$ and under the stated supposition of the theorem, we have for any $\tilde{q}(x,{\mathcal{F}}_{p^{\ast}})\in\tilde{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi)$ and any $x\in{{\mathbb{X}}}$
$\tilde{q}(x,{\mathcal{F}}_{p^{\ast}}):=q(x,{\mathcal{F}}_{p^{\ast}}(x))=q(x,\pi^{\ast}(x))=:\tilde{q}(x,\pi^{\ast}),$ (11)
noting that $p^{\ast}$ is the unique vector in the parameter set ${{\mathbb{S}}}$ associated with $\pi^{\ast}(x)$. Let ${q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}(x))=\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ denote the action-value function that corresponds by definition to $\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ of (11) in accordance with (10). Then, in one direction to show $\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})\leq q_{B}^{\ast}(x,\pi^{\ast}(x))$, we derive
$q_{B}^{\ast}(x,\pi^{\ast}(x))$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,\pi^{\ast})\Big{[}r(x,\pi^{\ast}(x))+\gamma q_{B}^{\ast}(y,\pi^{\ast}(y))\Big{]}$
$\stackrel{{\scriptstyle(b)}}{{\geq}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,\pi^{\ast})\Big{[}r(x,\pi^{\ast}(x))+\gamma{q}_{C}^{\ast}(y,\pi^{\ast}(y))\Big{]}\stackrel{{\scriptstyle(c)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\ast}}(y))\Big{]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma\tilde{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\ast}})\Big{]}\stackrel{{\scriptstyle(e)}}{{=}}\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}),$ (12)
where: (a) is directly from the definition of the fixed point in (9); (b) follows from the definition of the Bellman operator ${{\mathcal{T}}_{B}}$ for any action-value function $q(x,u)\in{\mathbb{Q}}({{\mathbb{X}}}\times{\mathbb{U}})$, the unique fixed point (9), and the optimality of $q_{B}^{\ast}$ and $\pi^{\ast}$; (c) follows upon the substitution of policy ${\mathcal{F}}_{p^{\ast}}$ for $\pi^{\ast}$ and ${q}_{C}^{\ast}(x,\pi^{\ast}(x))={q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}(x))$ from (11); (d) follows upon the substitution of ${q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}(x))=\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ from (11); and (e) follows from (10). In the other direction to show $q_{B}^{\ast}(x,\pi^{\ast}(x))\leq\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$, we derive
$\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma\tilde{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\ast}})\Big{]}\stackrel{{\scriptstyle(b)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma{q}_{C}^{\ast}(y,{\mathcal{F}}_{p^{\ast}}(y))\Big{]}$
$\stackrel{{\scriptstyle(c)}}{{\geq}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p^{\ast}})\Big{[}r(x,{\mathcal{F}}_{p^{\ast}}(x))+\gamma{q}_{B}^{\ast}(y,{\mathcal{F}}_{p^{\ast}}(y))\Big{]}\stackrel{{\scriptstyle(d)}}{{=}}\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,\pi^{\ast})\Big{[}r(x,\pi^{\ast}(x))+\gamma{q}_{B}^{\ast}(y,\pi^{\ast}(y))\Big{]}$
=$\stackrel{{\scriptstyle(e)}}{{=}}q_{B}^{\ast}(x,\pi^{\ast}(x)),$ (13)
where: (a) is directly from the definition of the fixed point in (10); (b) follows upon the substitution of ${q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}(x))=\tilde{q}_{C}^{\ast}(x,{\mathcal{F}}_{p^{\ast}})$ from (11); (c) follows from the definition of the CBRL operator ${{\mathbf{T}}}$ for any action-value function $q(x,u)\in{\mathbb{Q}}({{\mathbb{X}}}\times{\mathbb{U}})$, the unique fixed point (10), and the application of the Bellman optimality conditions on $\tilde{q}_{C}^{\ast}$ in (10) for ${\mathcal{F}}_{p^{\ast}}$, all under the supposition of the theorem; (d) follows upon the substitution of policy $\pi^{\ast}$ for ${\mathcal{F}}_{p^{\ast}}$ and $q_{B}^{\ast}(x,{\mathcal{F}}_{p^{\ast}}(x))=q_{B}^{\ast}(x,\pi^{\ast}(x))$ from (11); and (e) follows from (9). We therefore have coincidence of the fixed points (9) and (10), thus completing the proof.
Proof of Theorem 2.
@@
Suppose ${\mathbb{F}}$ satisfies Assumption 1, and consider a sequence of less rich families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k}\subset{\mathbb{F}}$ of policy functions ${\mathcal{G}}^{(i)}\in{\mathbb{F}}_{i}$ derived from parameter vectors of the corresponding parameter sets ${{\mathbb{S}}}_{i}$, further defining the operators ${{\mathbf{T}}}_{i}:\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$ as in (CBRL Approach) for any function $\tilde{q}_{i}(x,{\mathcal{F}}_{p}^{(i)})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$, $i\in[k]:=\{1,\ldots,k\}$. From Theorem 1, for $i\in[k]$, we have that the contraction operators ${{\mathbf{T}}}_{i}:{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow{{\mathbb{Q}}}({{\mathbb{X}}}\times\Pi_{i})$ under the parameter sets ${{\mathbb{S}}}_{i}$ converge to the unique fixed points $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ which satisfy, for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$,
$\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}^{(i)}_{p})=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}^{(i)}_{p})\bigg{[}r(x,{\mathcal{F}}^{(i)}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}_{i}}\tilde{q}_{i}^{\prime}(y,{\mathcal{F}}^{(i)}_{p^{\prime}})\bigg{]}$ (14)
@@
corresponding to (4). Let us first consider two such families ${\mathbb{F}}_{1}\subset{\mathbb{F}}_{2}$, for which we introduce the following lemma used in the proof of Theorem 2.
Lemma 1.